# LIA: Supervised Fine-Tuning of Large Language Models for Automatic Issue Assignment

### v1

@@ 1. Introduction
In software development, issue tracking systems such as Jira, Bugzilla, and GitHub are used to manage and resolve issues reported by users or contributors. When a new issue is submitted, it is typically reviewed by a project maintainer or issue assigner, who verifies its validity, determines its severity, and assigns it to an appropriate developer (wu2022spatial; zhou2025issuecourier). This process, known as issue assignment, is essential for ensuring timely and accurate issue resolution. When issue assignment is ineffective, issues are often reassigned multiple times before reaching the right developer. For instance, in the Eclipse project, issue# 16036 (eclipsebug16036; eclipsebug16036activity) was reassigned more than a dozen times and took nearly 100 days to resolve, illustrating how manual assignment can be inefficient, increase maintenance costs, and delay software delivery (yadav2024developer; samir2023interpretable; wu2022spatial).
@@
Effective issue assignment is a challenging task. It requires detailed knowledge of the codebase, project structure, and the expertise and availability of contributors (hajari2024sofiawl; aini2025expertise). These challenges are even more significant in open-source projects, where contributor roles are flexible and developer expertise is not always well documented. As a result, manual assignment can become inconsistent and error-prone (zhou2025issuecourier). Furthermore, as software projects grow, the number of issue reports increases rapidly. Large-scale open-source projects like Eclipse, Mozilla, and Google Chromium receive thousands of new issue reports each month (xie2021devrec; zhou2025issuecourier), making manual assignment increasingly unsustainable. This has led to the development of automated issue assignment solutions, which generally fall into two categories: text-based and graph-based methods (dong2024neighborhood).
@@
Text-based approaches frame issue assignment as a supervised learning task, where machine learning (ML) and deep learning (DL) techniques are used to predict the most suitable developer based on textual information such as issue titles and descriptions (dipongkor2023comparative; lee2022light; arnob2025bug; aung2022multi; mikolov2013efficient; mani2019deeptriage; wang2024empirical). However, these methods come with notable limitations. They typically require large amounts of labeled data to perform well, which makes them less practical for smaller projects with limited issue histories (zhou2025issuecourier; arnob2025bug; akhavan2025linkanchorautonomousllmbasedagent; zhang2023ealink). Such dependence on extensive labeled data reduces the generalizability of text-based methods and limits their effectiveness in real-world settings.
@@
To overcome some of these challenges, state-of-the-art approaches have explored graph-based methods that represent relationships between developers, issue reports, and software artifacts (zhou2025issuecourier; cao2025complex; wu2022spatial; tao2025structural; dong2024neighborhood; dai2023graph). These approaches often use graph neural networks (GNNs) (wu2020comprehensive) to learn embeddings that reflect structural, semantic, and temporal information. This helps model patterns such as developers fixing issues similar to ones they resolved in the past (wu2022spatial). While graph-based methods can be effective, they are typically computationally expensive and rely on techniques like graph sampling or random walks (wu2022spatial), which may miss important connections. They also require explicit links between issues and developers, which are often incomplete or noisy (dong2024neighborhood).
@@
Recent advancements in large language models (LLMs) have shown that they can achieve outstanding results on a variety of software maintenance tasks (aracena2025applying; izadi2022predicting; nikeghbal2024girt; xia2023automated; sarafraz2024domain; akhavan2025linkanchorautonomousllmbasedagent; Izadi:EMSE2021; yadollahi2025enhancing; abedini2025hybrid; Abedini:TOSEM2024). This success is largely attributed to LLMs’ pre-training on massive and diverse corpora, including software-related text and source code, which enables them to learn both general linguistic patterns and domain-specific programming knowledge (aracena2025applying). As a result, LLMs offer a promising path to overcome two key limitations of prior issue assignment approaches: (i) the dependence of text-based methods on large labeled datasets, and (ii) the reliance of graph-based methods on complete and noise-free relational data, which is often impractical to construct in real-world settings. Furthermore, recent studies have shown that LLMs demonstrate stronger results on software maintenance tasks when they are domain-adapted through fine-tuning or instruction tuning (jin2023inferfix; aracena2025applying; zhang2023ealink; huang2025comprehensive; li2024exploratory; heo2025study).
@@
Motivated by these insights, We introduce an LLM-based Issue Assignment technique, called LIA. We fine-tune an open-source LLM model, i.e., DeepSeek-R1-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability), using supervised fine-tuning (radford2018improving; ouyang2022training; zhang2025instructiontuninglargelanguage) to help it learn the relationship between the issue report text and the developers who resolved them. We fine-tune the LLM for issue assignment using historical issue-assignee pairs. Each training instance consists of an issue title and description paired with its corresponding assignee. Through this process, the model learns associations between assignee identifiers and the semantic features of the issues they typically resolve. At inference time, when presented with unseen issues, the model leverages these learned patterns to generate a ranked list of candidate assignees.
@@
We applied this fine-tuning to two widely used datasets in issue assignment research: EclipseJDT and Mozilla (bettenburg2008duplicate). These datasets contain thousands of real-world issue reports along with the developers who fixed them. By learning from these examples, LIA can effectively recommend likely developers for new issue reports based solely on the issue text.
@@
To evaluate the effectiveness of LIA, we conduct a comprehensive study comparing our fine-tuned model to both its pretrained base (DeepSeek-R1-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability)) and to state-of-the-art issue assignment techniques. Specifically, we benchmark LIA and the base LLM against four representative approaches: NCGBT (dong2024neighborhood), GCBT (dai2023graph), GRCNN (wu2022spatial), and CBR (anvik2006should). We design our evaluation to answer two main research questions:
@@
  • RQ1: How does LIA perform compared to the original pretrained model on the task of issue assignment? LIA consistently outperformed the base model under zero-shot evaluation conditions across both datasets. The fine-tuned model achieved substantial gains in issue assignment accuracy, with improvements of up to +187.8% at Hit@1 on EclipseJDT and +6.3% at Hit@7 on Mozilla over its general-purpose pretrained base (i.e., DeepSeek-R1-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability)).
@@ 2. Motivation
Manual issue assignment often introduces inefficiencies and delays in the software maintenance process. One particularly illustrative case is Eclipse issue 16036 (eclipsebug16036; eclipsebug16036activity), titled “UI not responsive when stack is deep”. It shows a complex and extended reassignment pattern that exemplifies the practical challenges of manual developer assignment.
@@
A similar case can be found in Mozilla issue 33654 (mozilla33654), titled “TEXTAREA incorrectly applying ROWS= and COLS= (horizontal / vertical scrollbar extra space, with overlay scrollbars disabled)”, whose lifespan is visualized in Figure 1. This issue remained open for nearly 25 years, offering a striking example of how inefficient assignment can delay even straightforward fixes. The issue was first reported in 1999 and immediately passed through multiple reassignments as project maintainers struggled to determine appropriate ownership. It was initially assigned and then reassigned several times in the following year, reflecting the uncertainty about which developer had the right expertise to address the problem. Despite early attempts at resolution, including patches that led to the issue being marked as resolved, it was reopened multiple times when the issue resurfaced in testing. The report continued to change hands, with six different developers becoming involved over time. At several points, it was completely unassigned, reverting to Mozilla’s default owner due to a lack of clarity or commitment. Even developers who began implementing partial fixes ultimately unassigned themselves, further delaying progress. Over the next two decades, the issue saw little forward momentum. Its status shifted repeatedly between resolved and reopened, with no final fix accepted. This iterative cycle of reassignment and reopening exemplifies the challenges of manual issue assignment in large projects, where the absence of automated support and clear ownership can lead to persistent stagnation. In 2023, nearly after 25 years of its original filing, the issue was finally closed by marking it as a duplicate of a newer report that had implemented the necessary correction.
These cases highlight how manual assignment introduces significant inefficiencies into the issue assignment process. When issues are assigned to wrong developers, they often remain unresolved, resulting in unnecessary transfers and frequent reopenings. These delays, caused by unclear ownership and ineffective coordination, extend the time to resolution. Our work introduces LIA, a supervised fine-tuned (radford2018improving; ouyang2022training; zhang2025instructiontuninglargelanguage) LLM designed to automate issue assignment based on the textual content of issue reports. By training the model on historical issue data, LIA reduces the likelihood of initial misassignment and helps streamline the assignment process. This is especially valuable in large, collaborative projects where manual assignment may be inconsistent or overloaded.
@@ 3. Approach
In this section, we present a comprehensive overview of our methodology for adapting a pretrained LLM (DeepSeek-R1-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability)) to the task of automated issue assignment. As illustrated in Figure 2, our approach consists of two main stages: (1) data preparation and preprocessing and (2) supervised fine-tuning (radford2018improving; ouyang2022training; zhang2025instructiontuninglargelanguage) with instructional prompts. We begin by collecting and filtering issue tracker datasets from two large-scale open-source projects, EclipseJDT and Mozilla (bettenburg2008duplicate). These reports are chronologically sorted, split into training and test sets, and formatted into instruction-style prompts. The model is fine-tuned using causal language modeling loss, learning to predict the most appropriate developer based solely on issue text. Through this process, the model learns from historical issue-to-developer assignments and uses those patterns to identify developers most likely to handle new issues based on similarity to past tasks.
@@ 3.1. Datasets
We conducted our study using two large-scale datasets from established open-source software ecosystems, namely EclipseJDT and Mozilla (bettenburg2008duplicate). Both projects maintain extensive public issue trackers, which made them ideal candidates for studying automated issue assignment in a practical and reproducible setting. These datasets have been widely used in prior work (dong2024neighborhood) and offer rich, mature histories of issue reports and assignee annotations, as illustrated in Figure 3.
@@ 3.2. Preprocessing
To prevent temporal leakage, we adopted the same chronological partitioning strategy as NCGBT (dong2024neighborhood), which was evaluated on the same datasets. Specifically, for each dataset, issue reports were sorted by their creation timestamps. Each dataset was partitioned into 80% training, 10% validation, and 10% test portions. To ensure robustness and reduce potential bias from data ordering, we conducted multiple test runs in which the data was reshuffled and a different 10% subset was used for evaluation in each run. The reported results represent the averaged performance across these runs, providing a more stable and reliable measure of the model’s effectiveness.
Following NCGBT (dong2024neighborhood), to ensure data quality and reduce label noise, we applied a filtering step that removed developers who had contributed to fewer than 10 resolved issues. All issue reports associated with these infrequent developers were excluded from the dataset. This threshold helped focus the learning process on consistently active contributors and reduced the likelihood of overfitting to sparsely represented classes.
@@ 3.3. Supervised Fine-Tuning with Instructional Prompts
Supervised fine-tuning (SFT) (radford2018improving; ouyang2022training; zhang2025instructiontuninglargelanguage) is an effective method for adapting pretrained language models to specific tasks. Unlike general pretraining, where models learn broad language patterns from large amounts of unlabeled text, SFT focuses on teaching the model how to respond to specific types of inputs by training it on pairs of inputs and known outputs. In this setup, the model learns to generate correct responses when given structured instructions or queries, guided by examples that demonstrate the desired behavior.
In our project, we applied SFT to adapt the DeepSeek-R1-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability) model for the task of issue assignment in software issue assignment. The goal was to train the model to predict the most appropriate developer to handle a given issue, using only its textual description. To do this, we fine-tuned the model on historical issue reports from EclipseJDT and Mozilla (bettenburg2008duplicate) datasets. They contain thousands of resolved issues annotated with their actual fixers.
@@ 3.3.3. Training Configuration
  • Optimizer: AdamW (loshchilov2019decoupled) with weight decay set to 0.01; fused implementation used when supported.
@@ 4.1. Evaluation Metric
Following prior work (dong2024neighborhood), we evaluate with Hit@K, the proportion of test cases in which the true developer appears in the top-$K$:
@@ 4.3.1. Baseline for RQ1
We use the original DeepSeek-R1-Distill-Llama-8B model (deepseekai2025deepseekr1incentivizingreasoningcapability) as our baseline. This model, trained on general web text without any task-specific fine-tuning, serves as a strong zero-shot reference point to assess the effectiveness of domain adaptation for issue assignment.
@@ 4.3.2. Baselines for RQ2
  • NCGBT (dong2024neighborhood): A graph-based model designed for issue assignment that uses graph neural networks to model relationships from issue reports and developer collaboration networks for improved recommendations.
  • GCBT (dai2023graph): A graph-based collaborative filtering model that constructs a heterogeneous graph of issue and developers. It initializes issue nodes using a GRU-based text encoder, then applies a graph neural network to learn representations of both issues and developers. Predictions are made using an information retrieval strategy.
  • GRCNN (wu2022spatial): This model uses an LSTM-based module for issue reports and separately models developer relationships as a graph. It applies random walks to generate developer sequences from the graph and combines them with the textual information for final prediction.
  • CBR (anvik2006should): This text-based method uses a bag of words based on word frequency to convert the issue description into a vector, and then classifies it with SVM.

### v2

# arXiv:2601.01780v2[cs.SE] 27 Feb 2026

@@ 1. Introduction
In software development, issue tracking systems such as Jira, Bugzilla, and GitHub are used to manage and resolve issues reported by users or contributors. When a new issue is submitted, it is typically reviewed by a project maintainer or issue assigner, who verifies its validity, determines its severity, and assigns it to an appropriate developer (Wu et al., 2022; Zhou et al., 2025). This process, known as issue assignment, is essential for ensuring timely and accurate issue resolution. When issue assignment is ineffective, issues are often reassigned multiple times before reaching the right developer. For instance, in the Eclipse project, issue# 16036 (Eclipse Foundation, 2002a, b) was reassigned more than a dozen times and took nearly 100 days to resolve, illustrating how manual assignment can be inefficient, increase maintenance costs, and delay software delivery (Yadav et al., 2024; Samir et al., 2023; Wu et al., 2022).
@@
Effective issue assignment is a challenging task. It requires detailed knowledge of the codebase, project structure, and the expertise and availability of contributors (Hajari et al., 2024; Aini and Yulianti, 2025). These challenges are even more significant in open-source projects, where contributor roles are flexible and developer expertise is not always well documented. As a result, manual assignment can become inconsistent and error-prone (Zhou et al., 2025). Furthermore, as software projects grow, the number of issue reports increases rapidly. Large-scale open-source projects like Eclipse, Mozilla, and Google Chromium receive thousands of new issue reports each month (Xie et al., 2021; Zhou et al., 2025), making manual assignment increasingly unsustainable. This has led to the development of automated issue assignment solutions, which generally fall into two categories: text-based and graph-based methods (Dong et al., 2024).
@@
Text-based approaches frame issue assignment as a supervised learning task, where machine learning (ML) and deep learning (DL) techniques are used to predict the most suitable developer based on textual information such as issue titles and descriptions (Dipongkor and Moran, 2023; Lee et al., 2022; Arnob et al., 2025; Aung et al., 2022; Mikolov, 2013; Mani et al., 2019; Wang et al., 2024). However, these methods come with notable limitations. They typically require large amounts of labeled data to perform well, which makes them less practical for smaller projects with limited issue histories (Zhou et al., 2025; Arnob et al., 2025; Akhavan et al., 2025; Zhang et al., 2023). Such dependence on extensive labeled data reduces the generalizability of text-based methods and limits their effectiveness in real-world settings.
@@
To overcome some of these challenges, state-of-the-art approaches have explored graph-based methods that represent relationships between developers, issue reports, and software artifacts (Zhou et al., 2025; Cao and Cui, 2025; Wu et al., 2022; Tao et al., 2025; Dong et al., 2024; Dai et al., 2023). These approaches often use graph neural networks (GNNs) (Wu et al., 2020) to learn embeddings that reflect structural, semantic, and temporal information. This helps model patterns such as developers fixing issues similar to ones they resolved in the past (Wu et al., 2022). While graph-based methods can be effective, they are typically computationally expensive and rely on techniques like graph sampling or random walks (Wu et al., 2022), which may miss important connections. They also require explicit links between issues and developers, which are often incomplete or noisy (Dong et al., 2024).
@@
Recent advancements in large language models (LLMs) have shown that they can achieve outstanding results on a variety of software maintenance tasks (Aracena et al., 2025; Izadi et al., 2022; Nikeghbal et al., 2024; Xia et al., 2023; Sarafraz et al., 2024; Akhavan et al., 2025; Izadi et al., 2021; Yadollahi et al., 2025; Abedini and Heydarnoori, 2025, 2024). This success is largely attributed to LLMs’ pre-training on massive and diverse corpora, including software-related text and source code, which enables them to learn both general linguistic patterns and domain-specific programming knowledge (Aracena et al., 2025). As a result, LLMs offer a promising path to overcome two key limitations of prior issue assignment approaches: (i) the dependence of text-based methods on large labeled datasets, and (ii) the reliance of graph-based methods on complete and noise-free relational data, which is often impractical to construct in real-world settings. Furthermore, recent studies have shown that LLMs demonstrate stronger results on software maintenance tasks when they are domain-adapted through fine-tuning or instruction tuning (Jin et al., 2023; Aracena et al., 2025; Zhang et al., 2023; Huang et al., 2025b; Li et al., 2024; Heo and Lee, 2025).
@@
Motivated by these insights, We introduce an LLM-based Issue Assignment technique, called LIA. We fine-tune an open-source LLM model, i.e., DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025), using supervised fine-tuning (Radford et al., 2018; Ouyang et al., 2022; Zhang et al., 2025) to help it learn the relationship between the issue report text and the developers who resolved them. We fine-tune the LLM for issue assignment using historical issue-assignee pairs. Each training instance consists of an issue title and description paired with its corresponding assignee. Through this process, the model learns associations between assignee identifiers and the semantic features of the issues they typically resolve. At inference time, when presented with unseen issues, the model leverages these learned patterns to generate a ranked list of candidate assignees.
@@
We applied this fine-tuning to two widely used datasets in issue assignment research: EclipseJDT and Mozilla (Bettenburg et al., 2008). These datasets contain thousands of real-world issue reports along with the developers who fixed them. By learning from these examples, LIA can effectively recommend likely developers for new issue reports based solely on the issue text.
@@
To evaluate the effectiveness of LIA, we conduct a comprehensive study comparing our fine-tuned model to both its pretrained base (DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025)) and to state-of-the-art issue assignment techniques. Specifically, we benchmark LIA and the base LLM against four representative approaches: NCGBT (Dong et al., 2024), GCBT (Dai et al., 2023), GRCNN (Wu et al., 2022), and CBR (Anvik et al., 2006). We design our evaluation to answer two main research questions:
@@
  • RQ1: How does LIA perform compared to the original pretrained model on the task of issue assignment? LIA consistently outperformed the base model under zero-shot evaluation conditions across both datasets. The fine-tuned model achieved substantial gains in issue assignment accuracy, with improvements of up to +187.8% at Hit@1 on EclipseJDT and +6.3% at Hit@7 on Mozilla over its general-purpose pretrained base (i.e., DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025)).
@@ 2. Motivation
Manual issue assignment often introduces inefficiencies and delays in the software maintenance process. One particularly illustrative case is Eclipse issue 16036 (Eclipse Foundation, 2002a, b), titled “UI not responsive when stack is deep”. It shows a complex and extended reassignment pattern that exemplifies the practical challenges of manual developer assignment.
@@
A similar case can be found in Mozilla issue 33654 (Mozilla Bugzilla, ), titled “TEXTAREA incorrectly applying ROWS= and COLS= (horizontal / vertical scrollbar extra space, with overlay scrollbars disabled)”, whose lifespan is visualized in Figure 1. This issue remained open for nearly 25 years, offering a striking example of how inefficient assignment can delay even straightforward fixes. The issue was first reported in 1999 and immediately passed through multiple reassignments as project maintainers struggled to determine appropriate ownership. It was initially assigned and then reassigned several times in the following year, reflecting the uncertainty about which developer had the right expertise to address the problem. Despite early attempts at resolution, including patches that led to the issue being marked as resolved, it was reopened multiple times when the issue resurfaced in testing. The report continued to change hands, with six different developers becoming involved over time. At several points, it was completely unassigned, reverting to Mozilla’s default owner due to a lack of clarity or commitment. Even developers who began implementing partial fixes ultimately unassigned themselves, further delaying progress. Over the next two decades, the issue saw little forward momentum. Its status shifted repeatedly between resolved and reopened, with no final fix accepted. This iterative cycle of reassignment and reopening exemplifies the challenges of manual issue assignment in large projects, where the absence of automated support and clear ownership can lead to persistent stagnation. In 2023, nearly after 25 years of its original filing, the issue was finally closed by marking it as a duplicate of a newer report that had implemented the necessary correction.
These cases highlight how manual assignment introduces significant inefficiencies into the issue assignment process. When issues are assigned to wrong developers, they often remain unresolved, resulting in unnecessary transfers and frequent reopenings. These delays, caused by unclear ownership and ineffective coordination, extend the time to resolution. Our work introduces LIA, a supervised fine-tuned (Radford et al., 2018; Ouyang et al., 2022; Zhang et al., 2025) LLM designed to automate issue assignment based on the textual content of issue reports. By training the model on historical issue data, LIA reduces the likelihood of initial misassignment and helps streamline the assignment process. This is especially valuable in large, collaborative projects where manual assignment may be inconsistent or overloaded.
@@ 3. Approach
In this section, we present a comprehensive overview of our methodology for adapting a pretrained LLM (DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025)) to the task of automated issue assignment. As illustrated in Figure 2, our approach consists of two main stages: (1) data preparation and preprocessing and (2) supervised fine-tuning (Radford et al., 2018; Ouyang et al., 2022; Zhang et al., 2025) with instructional prompts. We begin by collecting and filtering issue tracker datasets from two large-scale open-source projects, EclipseJDT and Mozilla (Bettenburg et al., 2008). These reports are chronologically sorted, split into training and test sets, and formatted into instruction-style prompts. The model is fine-tuned using causal language modeling loss, learning to predict the most appropriate developer based solely on issue text. Through this process, the model learns from historical issue-to-developer assignments and uses those patterns to identify developers most likely to handle new issues based on similarity to past tasks.
@@ 3.1. Datasets
We conducted our study using two large-scale datasets from established open-source software ecosystems, namely EclipseJDT and Mozilla (Bettenburg et al., 2008). Both projects maintain extensive public issue trackers, which made them ideal candidates for studying automated issue assignment in a practical and reproducible setting. These datasets have been widely used in prior work (Dong et al., 2024) and offer rich, mature histories of issue reports and assignee annotations, as illustrated in Figure 3.
@@ 3.2. Preprocessing
To prevent temporal leakage, we adopted the same chronological partitioning strategy as NCGBT (Dong et al., 2024), which was evaluated on the same datasets. Specifically, for each dataset, issue reports were sorted by their creation timestamps. Each dataset was partitioned into 80% training, 10% validation, and 10% test portions. To ensure robustness and reduce potential bias from data ordering, we conducted multiple test runs in which the data was reshuffled and a different 10% subset was used for evaluation in each run. The reported results represent the averaged performance across these runs, providing a more stable and reliable measure of the model’s effectiveness.
Following NCGBT (Dong et al., 2024), to ensure data quality and reduce label noise, we applied a filtering step that removed developers who had contributed to fewer than 10 resolved issues. All issue reports associated with these infrequent developers were excluded from the dataset. This threshold helped focus the learning process on consistently active contributors and reduced the likelihood of overfitting to sparsely represented classes.
@@ 3.3. Supervised Fine-Tuning with Instructional Prompts
Supervised fine-tuning (SFT) (Radford et al., 2018; Ouyang et al., 2022; Zhang et al., 2025) is an effective method for adapting pretrained language models to specific tasks. Unlike general pretraining, where models learn broad language patterns from large amounts of unlabeled text, SFT focuses on teaching the model how to respond to specific types of inputs by training it on pairs of inputs and known outputs. In this setup, the model learns to generate correct responses when given structured instructions or queries, guided by examples that demonstrate the desired behavior.
In our project, we applied SFT to adapt the DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025) model for the task of issue assignment in software issue assignment. The goal was to train the model to predict the most appropriate developer to handle a given issue, using only its textual description. To do this, we fine-tuned the model on historical issue reports from EclipseJDT and Mozilla (Bettenburg et al., 2008) datasets. They contain thousands of resolved issues annotated with their actual fixers.
@@ 3.3.3. Training Configuration
  • Optimizer: AdamW (Loshchilov and Hutter, 2019) with weight decay set to 0.01; fused implementation used when supported.
@@ 4.1. Evaluation Metric
Following prior work (Dong et al., 2024), we evaluate with Hit@K, the proportion of test cases in which the true developer appears in the top-$K$:
@@ 4.3.1. Baseline for RQ1
We use the original DeepSeek-R1-Distill-Llama-8B model (DeepSeek-AI, 2025) as our baseline. This model, trained on general web text without any task-specific fine-tuning, serves as a strong zero-shot reference point to assess the effectiveness of domain adaptation for issue assignment.
@@ 4.3.2. Baselines for RQ2
  • NCGBT (Dong et al., 2024): A graph-based model designed for issue assignment that uses graph neural networks to model relationships from issue reports and developer collaboration networks for improved recommendations.
  • GCBT (Dai et al., 2023): A graph-based collaborative filtering model that constructs a heterogeneous graph of issue and developers. It initializes issue nodes using a GRU-based text encoder, then applies a graph neural network to learn representations of both issues and developers. Predictions are made using an information retrieval strategy.
  • GRCNN (Wu et al., 2022): This model uses an LSTM-based module for issue reports and separately models developer relationships as a graph. It applies random walks to generate developer sequences from the graph and combines them with the textual information for final prediction.
  • CBR (Anvik et al., 2006): This text-based method uses a bag of words based on word frequency to convert the issue description into a vector, and then classifies it with SVM.